Find the integral surface of quasi linear partial differential equation xzp+yzq=-xy which passes through the curve y=x²,z=x³



1
Expert's answer
2021-04-13T23:42:58-0400

"xzp+yzq=-xy
\\
\text{This is lagrange's equation}\\
Pp+Qq=R\\
\frac{dx}{p}=\frac{dy}{Q}=\frac{dz}{-xy}\\
\frac{dx}{xz}=\frac{dy}{yz} \implies \frac{dx}{x}=\frac{dy}{y}\\
\implies\int\frac{dx}{x}=\int\frac{dy}{y}\\
Inx=Iny+Ina\\
\implies In\frac{x}{y}=Ina\\
\frac{x}{y}=a\\
\frac{ydx+xdy-2zdz}{xyz+xyz-2xyz}=\frac{ydx+xdy-2zdz}{0}\\
d(xy)-2zdz=0\\
xy-z^2=b\\
\psi(\frac{x}{y},xy-z^2)=0."